# Understanding Self-Supervised Learning via Gaussian Mixture Models

### v2

@@ 4.1 Fisher Subspace
Given a SharedGMM (Def 3.1) parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$, its Fisher subspace Fukunaga (2013), denoted by ${S}_{F}$, is the smallest subspace that achieves the maximum Fisher discriminant, which is given by:
@@
Class Preserving Subspace.
Additionally, we use a unique property Fukunaga (2013) of the Fisher subspace; it is the smallest subspace preserving the class posterior probabilities for Gaussian mixtures with shared covariance.
@@
Crucially, this property implies that projecting a GMM onto its Fisher subspace will not result in mode collapse or erroneous mode merging. Moreover, any clustering algorithm operating in the lower dimensional subspace will observe the same class probabilities as in the original space. Owing to it’s properties of maximizing the ratio of inter to intra-component variances and preservation of class probabilities (i.e., index of underlying component), we argue that Fisher subspace is the optimal subspace for projection, particularly for the class of shared covariance GMMs.
4.2 Spectral Dimensionality Reduction : SVD-subspace
Spectral methods Kannan et al. (2009) are the standard methods for LDR. They are (generally) based on the principle of finding the best-fit subspace which maximally explains the data. This subspace is given by the top singular vectors of data covariance.
Definition 4.3.
The $r$-dimensional SVD subspace Vempala and Wang (2004) $S^{(r)}_{SVD}$ for a GMM $F$ is described as the top $r$ left singular vectors of $\mathbb{E}_{{{\bm{x}}}\sim F}[{{\bm{x}}}{{\bm{x}}}^{T}]$.
A vanilla spectral dimensionality reduction method hence projects the data points onto the $r$-dimensional SVD subspace.  Kannan et al. (2009) proved that for spherical GMMs (i.e. ${\bm{\Sigma}}={{\bm{I}}}$), SVD subspace is equal to the mean subspace $S_{{\bm{\mu}}}=\mathrm{Span}\big{(}{\bm{\mu}}\big{)}$. Moreover, using Eqn 1 we can derive that $S_{F}=S_{{\bm{\mu}}}=S_{SVD}$. Hence, projection onto the SVD subspace is the optimal dimensionality reduction method for spherical GMMs.
4.3 Parallel Pancakes
@@
Figure 1: (a) For spherical Gaussians, $S_{SVD}$ and $S_{F}$ overlap and hence projection onto $S_{SVD}$ leads to well seperated GMM. (b) (Parallel Pancakes) For general non-spherical Gaussians, large variance in some direction leads to $S_{SVD}\neq S_{F}$. Hence, projection onto $S_{SVD}$ leads to mode collapse. (c) (Shifted Parallel Pancakes) Mean subspace $S_{{\bm{\mu}}}$ does not always coincide with $S_{F}$.
@@
While SVD-subspace projection is known to be optimal for spherical GMMs, it is easy to construct examples where spectral methods fail. Intuitively, spectral methods aim to maximally explain the data. This objective leads to “bad” projections when the variance in certain directions is more than the separation between the means, for which we introduce the “parallel pancakes” example  Kannan et al. (2009). Consider a two component GMM that resembles “parallel pancakes”, i.e., the two Gaussians are narrow and separated along one direction, and spherical in all other directions (see Fig 1). The two dimensional SVD subspace for this GMM is given by a plane parallel to the pancakes. The Fisher subspace however, is the plane passing through the means; in formal terms $S_{\mathrm{SVD}}\not\subset S_{F}$. In fact, the SVD subspace has the smallest Fisher discriminant among all two-dimensional subspaces, and hence, is worse than projection onto a random plane. Finding the optimal low dimensional space for non-spherical GMMs is known to be a non-trivial Brubaker and Vempala (2008); Kannan et al. (2009).
5 Self-Supervised Learning in Gaussian Mixture Models
@@
We analyze solving the dimensionality reduction task for GMMs with two popular self-supervised methods, namely SimCLR Chen et al. (2020a) and SimSiam Chen and He (2021). Both methods build on using augmentation pairs, but differ in their optimization objective (specifically how they regularize). We show that both these (differing) objectives are able to effectively leverage the augmentations and learn mappings onto the (optimal) Fisher subspace.
5.1 Augmentation-enabled Distribution
We define the augmentation of a point ${{\bm{x}}}$ to be an independent sample from the mixture distribution with a bias towards the underlying component of ${{\bm{x}}}$. The bias implies that if ${{\bm{x}}}$ is sampled from a component $z$, then its augmentation, on average, is more likely to be sampled from the same component $z$. Mathematically, we have :
Definition 5.1 (Augmentation-enabled Distribution (AeD)).
For a SharedGMM $F$ parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$, we define its Augmentation-enabled Distribution  $\hat{F}$, parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}},\delta\}_{k\in[K]}$ as
$\hat{F}=\delta\underset{k\in[K]}{\sum}w_{k}\bigg{(}\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\times\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\bigg{)}$
$+(1-\delta)\bigg{(}\underset{k\in[K]}{\sum}w_{k}\mathcal{N}({\bm{\mu}}_{k},{\bm{\Sigma}})\times\underset{{k^{\prime}}\in[K]}{\sum}w_{k^{\prime}}\mathcal{N}({\bm{\mu}}_{k^{\prime}},{\bm{\Sigma}})\bigg{)}$
To elucidate, augmentation-enabled distribution (AeD) returns a pair of points $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}$in a two-step process. We first flip a coin with bias $\delta$. Based on the coin flip, we either sample twice from the same component or sample from a possibly different components (chosen with probability $w_{k}$). The bias of the coin $\delta$ controls the correlation between ${{\bm{x}}}$ and $\hat{{{\bm{x}}}})$. For $\delta=0$, ${{\bm{x}}}$ and $\hat{{{\bm{x}}}}$ are independent samples from $F$, while for $\delta=1$, ${{\bm{x}}}$ and $\hat{{{\bm{x}}}}$ are always from the same component. The definition ensures that the marginal of both ${{\bm{x}}}$ and $\hat{{{\bm{x}}}}$ are equal to $F$ for any $\delta$.
Connection to Linear Discriminant Analysis (LDA)
. LDA Fukunaga (2013) is a supervised LDR method that leverages the class information to learn the subspace where the data distribution conditioned on the class label (i.e. the underlying component’s index) is maximally separated (see Suppl H). We highlight that assuming AeD (Def. 5.1) is strictly weaker than assuming a labeled dataset. We show that this condition (i.e. access to Augmentation-enabled Distribution ) is enough for learning the Fisher subspace and gives similar performance to supervised LDR methods.
@@ 5.2 Optimization Objective
InfoNCE loss
Oord et al. (2018) is popularly used in contrastive learning methods (like SimCLR). InfoNCE loss aims to learn representations such that a point is close to its augmentation, but far apart from everything else in the representation space. Concretely, for linear mappings we have
$\begin{split}&\mathcal{L}_{\mathrm{Info}}({{\bm{A}}})=-\underset{({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}}{\mathbb{E}}\Bigg{[}({{\bm{A}}}^{T}{{\bm{x}}})^{T}{{\bm{A}}}^{T}\hat{{{\bm{x}}}}\Bigg{]}\\ &+\underset{{{\bm{x}}}\sim F}{\mathbb{E}}\Bigg{[}\log\bigg{(}\underset{\tilde{{{\bm{x}}}}\sim F}{\mathbb{E}}\bigg{[}\exp\big{(}({{\bm{A}}}^{T}{{\bm{x}}})^{T}{{\bm{A}}}^{T}\tilde{{{\bm{x}}}}\big{)}\bigg{]}\bigg{)}\Bigg{]}\end{split}$ (2)
@@
where $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}$ is a sample from AeD (Def 5.1). ${{\bm{x}}}\sim F$ and $\tilde{{{\bm{x}}}}\sim F$ are independent draws from the mixture distribution (equivalently $({{\bm{x}}},\cdot)\sim\hat{F}$). The attractive term (i.e. the first term) keeps ${{\bm{x}}}$ and $\hat{{{\bm{x}}}}$ close by maximizing their dot product, while regularization term penalizes “proximity” to random samples. Intuitively, the first term in InfoNCE loss aims to increase the inter-component variance while the second term aims to decrease data variance (and hence intra-component variance).
SimSiam Loss
@@
Chen and He (2021) is another popular self-supervised loss “without negatives”.
$\underset{({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}}{-\ \mathbb{E}}\Bigg{[}\bigg{(}\frac{{{\bm{A}}}_{p}^{T}{{\bm{A}}}^{T}{{\bm{x}}}}{||{{\bm{A}}}_{p}^{T}{{\bm{A}}}^{T}{{\bm{x}}}||_{2}}\bigg{)}^{T}\mathrm{StopGrad}\bigg{(}\frac{{{\bm{A}}}^{T}\hat{{{\bm{x}}}}}{||{{\bm{A}}}^{T}\hat{{{\bm{x}}}}||_{2}}\bigg{)}\Bigg{]}$
@@
Note that it only requires augmentations (from $\hat{F}$) and doesn’t use random samples/negatives (i.e., $\tilde{{{\bm{x}}}}\sim F$ in Eqn 2). The loss is parameterized by two matrices : ${{\bm{A}}}$ (the mapping) and ${{\bm{A}}}_{p}$ (an additional linear mapping not utilized for projection). The loss also includes a StopGrad operation, which zeros out the gradients flowing through it. For SimSiam, it is trivial to observe that it allows for a collapsing solution (i.e., mapping all points onto the same vector). Prior work Tian et al. (2021); Jing et al. (2021); Wen and Li (2022) has argued that ${{\bm{A}}}_{p}$ and StopGrad play a crucial role in the training dynamics, preventing the occurrence of collapse. Our goal is to analyze the fixed point of this loss function. To do so, we introduce some simplifying assumptions and examine a modified SimSiam loss:
$\mathcal{L}_{\mathrm{Siam}}({{\bm{A}}})=$ $-\underset{({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}}{\mathbb{E}}\Bigg{[}({{\bm{A}}}^{T}{{\bm{x}}})^{T}{{\bm{A}}}^{T}\hat{{{\bm{x}}}}\Bigg{]}$
$+\xi\underset{{{\bm{x}}}\sim F}{\mathbb{E}}\Bigg{[}||{{\bm{A}}}^{T}{{\bm{x}}}||^{2}\Bigg{]}$ (3)
@@
We remove the StopGrad operation and set the prediction head (i.e. ${{\bm{A}}}_{p}$) to ${{\bm{I}}}$. We also trade the normalization term with a regularization term weighted by $\xi$. The attractive term in SimSiam behaves similar to InfoNCE loss, while the regularization term penalizes the norm of the projected points. The loss scales linearly with the squared norm of the matrix ${{\bm{A}}}$; we impose a norm constraint when optimizing for ${{\bm{A}}}$.
Contrasting to best-fit objective
We have established that spectral methods (specifically, SVD-subspace projection) learn projection matrices that induces maximal variance in the projected space;
$\underset{\{{{\bm{A}}}\in\mathbb{R}^{d\times r}:{{\bm{A}}}^{T}{{\bm{A}}}=I\}}{\mathrm{argmax}}\underset{{{\bm{x}}}\sim F}{\mathbb{E}}[||{{\bm{A}}}^{T}{{\bm{x}}}||^{2}]$ (4)
Informally, this objective is orthogonal to the goal of dimensionality reduction. Recall that our dimensionality reduction definition aims to find a projection space such that intra-component variance is low while inter-component variance is high. We can observe that optimizing a self-supervised objective serves as a better proxy for dimensionality reduction. We will now formalize our claim.
@@ 5.3 Main theorem
Equipped with the definitions of augmentations (Def 5.1) and the optimization objectives (Eqn 25.2), we prove that we can find the optimal projection matrix for the class of SharedGMMs (Def. 3.1).
@@
${{\bm{A}}}_{\mathrm{Info}}=\underset{{{\bm{A}}}\in\mathbb{R}^{d\times r}}{\mathrm{argmin}}\ \mathcal{L}_{\mathrm{Info}}({{\bm{A}}}),\,\,\,{{\bm{A}}}_{\mathrm{Siam}}=\underset{\begin{subarray}{c}{{\bm{A}}}\in\mathbb{R}^{d\times r}\\ ||{{\bm{A}}}||_{2}\leq 1\\ \end{subarray}}{\mathrm{argmin}}\ \mathcal{L}_{\mathrm{Siam}}({{\bm{A}}})$
Then for some $r\geq K$, $S_{\mathrm{Info}}\triangleq\mathrm{Col}({{\bm{A}}}_{\mathrm{Info}})$, $S_{\mathrm{Siam}}\triangleq\mathrm{Col}({{\bm{A}}}_{\mathrm{Siam}})$, we have :
@@
Thm 5.2 states that access to Augmentation-enabled Distribution augmentations is provably sufficient for learning the Fisher subspace for SharedGMMs. The result implies that class labels — typically assumed by supervised LDR methods like LDA — are not necessary.
The proof of Thm 5.2 can be found in Suppl CD. For population setting, self-supervised methods learn mapping onto a subset of the Fisher space. This implies that the projected points do not capture any noise and only contain a subset of the useful features. Note that spectral methods do not have any guarantees of this form (See Sec 4.3 for discussion).
@@
The theorem also states if we have perfect augmentations for InfoNCE loss, i.e. $\delta=1$, we cover all directions in the Fisher space and learn it exactly. The equivalence is true for SimSiam loss when the regularization coefficient $\xi$ is upper bounded appropriately. Importantly, this precludes learning of the collapsing solution for SimSiam loss. Under these conditions, we learn the same subspace as completely supervised dimensionality reduction methods.
Technical Difficulties
@@
Although we do not provide a proof for $S_{\mathrm{Info}}=S_{F}$ when $10<\delta<$ 1, we conjecture that the same is true, for which we include an empirical discussion in Sec 7. The results in Fig 4(a) suggest that $S_{\mathrm{Info}}=S_{F}$ even for $\delta<1$, and we leave the concrete analysis in this value range as future work. Additionally, we do not have an exact characterization of InfoNCE solution (we characterize just the column space of ${{\bm{A}}}_{\mathrm{Info}}$). Our empirical results show that the mapped points are well-separated compared to a projection onto the Fisher subspace (see Fig 2(d)). Furthermore, while we prove that both contrastive and non-contrastive objectives learn the same subspace, further investigation is required for their direct comparison. In Suppl F we provide an example where ${{\bm{A}}}_{\mathrm{Info}}$ is strictly better than ${{\bm{A}}}_{\mathrm{SimSiam}}$ (for $r<K$). We also present an empirical study on the effect of $r$ in Section 7 (see Figure 4(c)).
6 Multi-Modal Gaussian Mixtures and Self-Supervised Learning

### v3

@@ 3 Fisher subspace
Given a SharedGMM (Def 2.1) parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$, its Fisher subspace [9], denoted by ${S}_{F}$, is the smallest subspace that achieves the maximum Fisher discriminant, which is given by:
@@
We also use a unique property of the Fisher subspace [9]; it is the smallest subspace preserving the class posterior probabilities for Gaussian mixtures with shared covariance, which we formalize next.
@@
Crucially, this property implies that projecting a GMM onto its Fisher subspace will not result in mode collapse or erroneous mode merging. Moreover, any clustering algorithm operating in the lower dimensional subspace will observe the same class probabilities as in the original space. Owing to these properties of the Fisher subspace, we argue that it is the optimal subspace for projection, particularly for the class of shared covariance GMMs. Under the fully labeled SharedGMM setting, one could deduce that the subspace learned by (multi-class) LDA, $S_{LDA}$, for SharedGMM coincides with Fisher subspace $S_{F}$.
3.1 A simple example: Parallel pancakes [17]
@@
Figure 1: (a) (Spherical Gaussians) $S_{SVD}=S_{LDA}$ and projection onto $S_{SVD}$ leads to well seperated GMM. (b) (Non-spherical Gaussians) Large variance in one direction means $S_{SVD}\neq S_{LDA}$ and projection onto $S_{SVD}$ leads to mode collapse. (c) (Shifted non-spherical Gaussians) $S_{{\bm{\mu}}}\neq S_{LDA}$.
Spectral methods [17] are the standard methods for LDR. They are based on the principle of finding the best-fit subspace which is (generally) given by the top singular vectors of data covariance. Formally, the $r$-dimensional SVD subspace [25] $S^{(r)}_{SVD}$ for a GMM $F$ is described as the top $r$ left singular vectors of $\mathbb{E}_{{{\bm{x}}}\sim F}[{{\bm{x}}}{{\bm{x}}}^{T}]$. A vanilla spectral dimensionality reduction method hence projects the data points onto the $r$-dimensional SVD subspace.  [17] proved that for spherical GMMs (i.e. ${\bm{\Sigma}}={{\bm{I}}}$), SVD subspace is equal to the mean subspace $S_{{\bm{\mu}}}=\mathrm{Span}\big({\bm{\mu}}\big)$. Moreover, using Eq. (1) we can derive that $S_{LDA}=S_{{\bm{\mu}}}=S_{SVD}$. Hence, projection onto the SVD subspace is the optimal dimensionality reduction method for spherical GMMs.
@@
However, it is easy to construct examples demanding for a sophisticated method. Spectral methods aim to maximally explain the data, leading to possibly “bad” projections when the variance in certain directions dominates the separation between the means. In Figure˜1 we present such an example; consider a two component GMM that resembles “parallel pancakes” such that the components are narrow and separated along one direction, and spherical in all other directions. The two dimensional SVD subspace for this GMM is given by a plane parallel to the pancakes. The Fisher subspace (and equivalently LDA subspace), however, is the plane passing through the means; formally, $S_{\mathrm{SVD}}\not\subset S_{F}$. In fact, the $S_{SVD}$ has the smallest Fisher discriminant among all two-dimensional subspaces. Finding the optimal subspace for non-spherical GMMs is known to be non-trivial [4, 17].
4 Contrastive learning for Gaussian mixture models
@@
We analyze solving the dimensionality reduction task for GMMs with two popular self-supervised methods, SimCLR and SimSiam. Both methods build on using augmentation pairs, but differ in their optimization objective (specifically how they regularize). We show that both objectives are able to effectively leverage the augmentations and learn mappings onto the (optimal) Fisher subspace.
@@ 4.1 Optimization objective
InfoNCE loss [19] is popularly used in contrastive learning methods (like SimCLR). It learns representations that pull data points and their augmentations close together while pushing them away from other points in the embedding space. We define the InfoNCE objective for linear mappings as,
[…]
@@
where $({{\bm{x}}},\hat{{{\bm{x}}}})\sim\hat{F}$ is a sample from augmentation-enabled distirbution (Def 2.2) such that ${{\bm{x}}}\sim F$ and $\tilde{{{\bm{x}}}}\sim F$ are independent draws from the mixture. The attractive term (i.e. the first term) keeps $\bm{x}$}} and $\hat{{{\bm{x}}}}$ close by maximizing their dot product, while regularization term penalizes proximity to random samples. In other terms, the first term increases inter-component variance while the second term decreases data variance and hence intra-component variance.
@@
SimSiam loss [7] is another popular self-supervised loss without the negatives.
[…]
@@
Unlike InfoNCE, Simsiam only operates on augmentations (from $\hat{F}$) and doesn’t use negatives. The loss is parameterized by two matrices; the mapping $\bm{A}$}} and the “prediction head” ${{\bm{A}}}_{p}$, which is not utilized for projection. Observe that Simsiam might be prone to a collapsing solution, i.e., mapping all points onto the same vector. Prior work [24, 15, 27] has argued that ${{\bm{A}}}_{p}$, which makes the optimization asymmetric, and StopGrad, which zeros out the gradients flowing through it, play a crucial role together in the guiding the training dynamics, preventing the occurrence of collapse. Our goal is to analyze the fixed point of this loss function. For tractability, we introduce some simplifying assumptions and examine a modified loss:
[…]
@@
We remove the StopGrad operation and set the prediction head (i.e. ${{\bm{A}}}_{p}$) to $\bm{I}$}}. We also trade the normalization term with a regularization term weighted by $\xi$. The attractive term in SimSiam behaves similar to InfoNCE loss, while the regularization term penalizes the norm of the projected points. The loss scales linearly with $\|{{\bm{A}}}\|^{2}$, therefore, we impose a norm constraint when optimizing for $\bm{A}$}}.
We have already established that maximizing total variance on its own in the projected space is not the preferable strategy for dimensionality reduction. In the presence of augmentation (and negatives), self-supervised objectives could take a finer-grained approach by maximizing inter-component variance and keeping intra-component variance low, simultaneously. Next, we will formalize our claim that self-supervised learning is a good proxy for fully supervised techniques (e.g., LDA).
@@ 4.2 Main theorem
Nex, we prove that InfoNCE and (simplified) Simsiam can find the Fisher-optimal projection matrix for the class of SharedGMMs.
@@
Suppose $F$ is a SharedGMM parameterized by $\{w_{k},{\bm{\mu}}_{k},{\bm{\Sigma}}\}_{k\in[K]}$ and $\hat{F}$ is its Augmentation-enabled Distribution with bias $\delta$. Let $S_{F}$ be the Fisher subspace (Eqn 1) of $F$. Denote $\lambda_{\mathrm{min}}$ as the minimum non-zero eigenvalue of ${\bm{\Sigma}}^{-\frac{1}{2}}\left(\sum_{k}w_{k}{\bm{\mu}}_{k}{\bm{\mu}}_{k}^{T}\right){\bm{\Sigma}}^{-\frac{1}{2}}$ and
@@
[…]
@@
Access to Augmentation-enabled Distribution is provably sufficient for learning the Fisher subspace for SharedGMMs matching the performance of supervised LDA. The result implies that class labels which are typically assumed by supervised LDR methods like LDA, are not necessary.
@@
The proof of Theorem˜4.1 can be found in Appendices˜C and D. For population setting, the fact that self-supervised methods learn mapping onto a subset of the Fisher space implies that the projected points do not capture any noise and only contain (a subset of) the useful features. Note that spectral methods do not have guarantees of this form. The theorem also states if we have perfect augmentations for InfoNCE loss, i.e. $\delta=1$, we cover all directions in the Fisher space and learn it exactly. The equivalence is true for SimSiam loss when the regularization coefficient $\xi$ is upper bounded appropriately. On a related front, this precludes learning of the collapsing solution for SimSiam loss, verifying the claims in prior work on a simplified version of the loss function. Most importantly, self-supervised methods in question learn the same subspace as supervised dimensionality reduction methods like LDA which needs the knowledge of underlying components for all the samples.
@@
Technical difficulties. Although we do not provide a proof for $S_{\mathrm{Info}}=S_{F}$ when $0<\delta<1$, we conjecture that it is true, for which we include an empirical discussion in Section˜6. Figure˜4(a) suggests that $S_{\mathrm{Info}}=S_{F}$ even for $\delta<1$, analysis of which we leave as future work. Additionally, we do not have an exact characterization of InfoNCE solution; we characterize just the column space of ${{\bm{A}}}_{\mathrm{Info}}$. Our empirical results show that the mapped points are well-separated compared to a projection onto the Fisher subspace (see Fig 2(d)). Furthermore, while we prove that both contrastive and non-contrastive objectives learn the same subspace, further investigation is required for their direct comparison. In Appendix˜F, we provide an example where ${{\bm{A}}}_{\mathrm{Info}}$ is strictly better than ${{\bm{A}}}_{\mathrm{SimSiam}}$ (for $r<K$). We also present an empirical study on the effect of $r$ (see Figure 4(c)).
5 Multi-modal Gaussian mixture models
Previously, we considered a setup where a point and its augmentation follow the same distribution. It is a natural assumption for methods like SimCLR and SimSiam where augmentations of points (i.e. images) are defined by transformations (e.g., cropping, color jittering) on the image. This assumption may not hold in general. Text embeddings models like DPR [18] or image-text embedding models like CLIP [21] consider input samples as a pair of points following (possibly) different distribution.